A jet is flying from Hawaii to San Fransisco a distance of 2400 miles. In still air the jet flies at 600mph but there is now a 40mph tailwind. In case of emergency how many hours after takeoff will it be faster for the jet to go on to San Francisco rather than return to Hawaii?
Answer by lwsshak3(11628)

A jet is flying from Hawaii to San Fransisco a distance of 2400 miles. In still air the jet flies at 600mph but there is now a 40mph tailwind. In case of emergency how many hours after takeoff will it be faster for the jet to go on to San Francisco rather than return to Hawaii?
**
let x=distance covered on the way to SF at turn-around point
2400-x=Remaining distance to SF at turn-around point
Speed of plane with tailwind=600+40=640mph (go on to SF)
Speed of plane against wind=600-40=560mph (rtn to Hawaii)
Travel time=distance/speed
Time to go on to SF =Time to return to Hawaii
(2400-x)/640=x/560
640x=560*2400-560x
1200x=560*2400
x=560*2400/1200=1120 miles
Travel time=(2400-1120)640=2hrs
ans:
After two hours after takeoff it will be faster for the jet to go on to San Francisco rather than return to Hawaii